# Benchmarking End-To-End Performance of AI-Based Chip Placement Algorithms

### v1

@@ 7.3 Discussion
Our benchmark comprises design kits essential for each stage of the physical design flow, including netlists, libraries, rules, and constraints needed during the physical implementation stage. This thorough inclusion allows for a convenient and detailed evaluation of algorithms at any specific stage of the physical design flow, enabling researchers and practitioners to test and compare the effectiveness of their solutions in a realistic, end-to-end environment. We call on the AI researchers to pay more attention on the “shift-left” challenge from the real-world industrial scenarios, keeping towards the mission of bridging the huge gap between academic research and industrial applications.
@@ 8 Conclusion
This paper presents a comprehensive dataset that spans the entire spectrum of the EDA design process and an end-to-end evaluation method, which we used to assess several placement algorithms: DREAMPlace, AutoDMP, MaskPlace, ChipFormer, Wiremask-EA, and SA. Our evaluation revealed inconsistencies between the metrics currently emphasized by mainstream placement algorithms and the final performance outcomes. These findings highlight the need for a new perspective in the development of placement algorithms.
@@ A.2 Encountered Errors
The code of ChiPFormer has certain issues and needs modifications to be applied to other cases. DREAMPlace directly uses the mix-size method with LEF/DEF for running macros.
For the ICCAD 2015 dataset, it includes the lib file, LEF/DEF files, netlist file, and sdc file. However, the lib file lacks buffer definitions, preventing CTS. The lef files have incomplete layer definitions, hindering routing.Additionaly, OpenRoad does not support the syntax of the ICCAD 2015 sdc file.
@@ Appendix C More Results
The detailed descriptions of each design are in Table 4. All the raw data from the experiment are in Tables 5-22. Methods or designs not shown in the table failed to complete the entire process due to the errors discussed earlier.
@@
1 CAN-BusTommydag A CAN bus Controller
3 FPGA-CANWangXuan95 A lightweight CAN bus controller
@@
8 black parrot The-OpenROAD-Project Full 64-bit RISC-V Core with Cache Coherence Directory
@@
10 camera interface AngeloJacobo Camera interface
11 image processing suntodai Image processing
@@
14 ariane136 The-OpenROAD-Project Ariane Core
16 chameleon The-OpenROAD-Project SoC core
17 riscv32i The-OpenROAD-Project Simple RISCV 32I core
@@
19 toygpu openrisc A simple GPU on a TinyFPGA BX
20 mor1kx matt-kimball an OpenRISC processor IP core
@@
WireMask-EA 1124169.895 5027796.1 0.227 6609025 0.369 -0.581 -687.560 2662 770528
SA 1683330.105 5109178.9 0.227 6594317 0.365 -0.812 -1368.340 3253 764028
DREAMPlace 1111023.474 4835183.2 0.222 6460013 0.370 -0.431 -578.455 2181 768426
AutoDMP 828592.368 4198935.5 0.195 5662040 0.351 -0.481 -622.147 2280 760842
MaskPlace 4444289.000 6258454.3 0.260 7720348 0.373 -0.847 -1553.500 3277 769790
ChiPFormer 1253799.737 4996747.8 0.229 6650789 0.372 -0.617 -933.311 2707 772172
OpenROAD 2685856.053 5226018.8 0.235 6817628 0.358 -0.492 -609.207 2279 770622
@@
WireMask-EA 1181404.421 5308713.2 0.243 7034760 0.578 -1.688 -4248.100 3546 791551
SA 1693986.737 5584824.9 0.252 7296870 0.574 -1.711 -4116.460 3662 792731
DREAMPlace 1067974.000 5157317.8 0.232 6885717 0.574 -1.380 -3293.090 3581 792608
AutoDMP 870515.895 4907336.9 0.229 6607637 0.569 -1.412 -3644.990 4071 791027
MaskPlace 4586233.421 6522052.1 0.290 8382804 0.586 -1.513 -3458.900 3608 796715
ChiPFormer 1269027.263 5078881.1 0.232 6705656 0.563 -1.348 -3266.360 3521 789667
OpenROAD 3067334.316 5540894.5 0.249 7176803 0.550 -1.318 -3204.870 3468 792446
Table 7: The results of black parrot
@@
WireMask-EA 112560.500 7310237.2 0.422 9637354 0.252 -1.827 -2.222 23 815781
SA 151696.447 6548499.6 0.385 8775996 0.248 -1.896 -165.720 925 813755
AutoDMP 135090.737 6338160.8 0.369 8427105 0.247 -1.951 -60.856 605 811220
DREAMPlace 161581.000 6798714.6 0.382 8914932 0.244 -1.884 -198.213 1398 813703
MaskPlace 324368.447 6808053.5 0.393 8966592 0.249 -1.657 -102.182 987 814013
ChiPFormer 138943.842 6893462.7 0.394 8993077 0.250 -1.746 -463.346 1687 816983
OpenROAD 119123.132 6242894.2 0.360 8209505 0.246 -1.961 -71.882 862 810360
Table 8: The results of bpfe
@@
WireMask-EA 45985.289 1566601.0 0.404 2206773 0.298 -1.013 -686.950 2263 234748
SA 59872.289 1289105.7 0.295 1633540 0.287 -1.060 -853.464 2459 237391
DREAMPlace 60280.000 1800046.2 0.498 2724715 0.301 -1.171 -808.973 2239 235832
AutoDMP 56107.000 1302271.3 0.293 1625619 0.292 -1.158 -765.161 2464 239894
MaskPlace 64024.000 1263243.8 0.285 1580587 0.285 -1.294 -641.347 1858 238130
ChiPFormer 45754.000 1282954.8 0.301 1670666 0.294 -0.967 -641.719 2457 235347
OpenROAD 53842.000 1167916.2 0.250 1389986 0.262 -1.007 -473.214 2441 231069
Table 9: The results of bpbe
@@
WireMask-EA 66729.237 1826669.2 0.331 2288717.000 0.397 -1.829 -2427.590 4816 262891.000
SA 85984.158 1939541.5 0.410 2816347.000 0.411 -1.787 -3412.900 5612 264921.000
DREAMPlace 87304.667 1777230.3 0.317 2219014.333 0.386 -1.772 -2294.093 4885 262545.667
AutoDMP 86263.842 1713273.4 0.307 2124910.000 0.384 -1.819 -2436.980 4909 261516.000
MaskPlace 114314.553 1909505.5 0.345 2386914.000 0.398 -1.778 -3458.640 4730 261446.000
ChiPFormer 62168.447 2175102.7 0.411 2830643.000 0.417 -1.803 -2611.130 5127 266110.000
OpenROAD 76561.237 1940820.8 0.337 2332089.000 0.389 -1.827 -2514.450 5034 259572.000
Table 10: The results of swerv wrapper
@@
WireMask-EA 474447.237 4186894.0 0.446 5209646 0.682 -1.127 -1256.030 1668 679048
SA 630919.500 4125523.0 0.455 5313519 0.682 -1.195 -880.450 1789 678647
DREAMPlace 865544.000 3673896.3 0.375 4631643 0.666 -0.889 -640.107 1165 675207
AutoDMP 450662.000 3171618.7 0.352 4130837 0.632 -0.711 -709.121 1495 672435
MaskPlace 1270726.500 4414508.8 0.461 5391952 0.681 -$1$743.094 1443 679852
ChiPFormer 437638.237 4056268.8 0.440 5144218 0.684 -1.079 -893.142 1756 679744
OpenROAD 666704.000 3187567.0 0.342 4015115 0.627 -0.978 -751.619 1531 673169
@@
Table 12: The results of FPGA-CAN
Method Intermediate Metrics PPA Metrics
AutoDMP 1437916.7 0.178 2393640 0.745 -0.892 -108.201 3864 379675
OpenROAD 1483498.8 0.180 2394544 0.742 -0.404 -30.331 1843 379193
Table 13: The results of CAN-Bus
Method Intermediate Metrics PPA Metrics
DREAMPlace 3916.8 0.093 5683 0.005 -0.074 -0.074 1 1675.270
AutoDMP 4210.8 0.097 6024 0.005 -0.074 -0.074 1 1668.880
OpenROAD 3955.1 0.092 5623 0.005 -0.079 -0.079 1 1665.430
@@
Table 16: The results of iot_shield
Method Intermediate Metrics PPA Metrics
DREAMPlace 5014.3 0.118 7346 0.006 -0.167 -2.457 27 1712.770
AutoDMP 5976.2 0.137 8653 0.006 -0.152 -2.367 29 1730.600
OpenROAD 5067.3 0.119 7349 0.006 -0.152 -2.186 20 1719.420
Table 17: The results of FPGA_image_processing
Method Intermediate Metrics PPA Metrics
DREAMPlace 16556.5 0.123 23453 0.201 -0.924 -184.592 736 5797.470
AutoDMP 22203.0 0.160 30745 0.211 -0.930 -200.125 710 5812.630
OpenROAD 16689.0 0.120 22746 0.199 -0.919 -165.009 609 5758.100
Table 18: The results of 8051-master
Method Intermediate Metrics PPA Metrics
DREAMPlace 142162.6 0.201 210127 0.075 -0.634 -16.460 35 29484.200
AutoDMP 131174.7 0.186 195536 0.074 -0.633 -15.643 35 29365.600
OpenROAD 141245.3 0.201 207737 0.073 -0.603 -16.664 41 29461.400
Table 19: The results of mor1kx
Method Intermediate Metrics PPA Metrics
Table 20: The results of toygpu
Method Intermediate Metrics PPA Metrics
DREAMPlace 4347733.8 0.198 6784639 1.101 -1.503 -131.661 101 979293
OpenROAD 4657703.6 0.212 7185236 1.108 -1.294 -103.668 101 981911
@@
AutoDMP 13446106.8 0.238 19114559.0 2.697 -0.67 -150.178 466 2319150.0
OpenROAD 14386152.9 0.253 20108941 3.425 -0.782 -305.819 674 2334330.0
Table 22: The results of Camera_Interface
Method Intermediate Metrics PPA Metrics

### v2

@@ 7.4 Discussion
Our benchmark comprises design kits essential for each stage of the physical design flow, including netlists, libraries, rules, and constraints needed during the physical implementation stage. This thorough inclusion allows for a convenient and detailed evaluation of algorithms at specific stages of the physical design flow, enabling researchers and practitioners to test and compare the effectiveness of their solutions in a realistic, end-to-end environment. We call on the AI researchers to pay more attention on the “shift-left” challenge from the real-world industrial scenarios, keeping towards the mission of bridging the huge gap between academic research and industrial applications.
@@ 8 Conclusion
This paper presents a comprehensive dataset that spans the entire spectrum of the EDA design process and an end-to-end evaluation method, which we used to assess several placement algorithms: SA, WireMask-EA, DREAMPlace, AutoDMP, MaskPlace, and ChiPFormer. Our evaluation revealed inconsistencies between the metrics currently emphasized by mainstream placement algorithms and the final performance outcomes. These findings highlight the need for a new perspective in the development of placement algorithms.
@@ A.2 Encountered Errors
The code of ChiPFormer has certain issues and needs modifications to be applied to other cases. DREAMPlace directly uses the mixed-size method with LEF/DEF for running macros.
For the ICCAD 2015 dataset, it includes the lib file, LEF/DEF files, netlist file, and sdc file. However, the lib file lacks buffer definitions, preventing CTS. The lef files have incomplete layer definitions, hindering routing.Additionally, OpenROAD does not support the syntax of the ICCAD 2015 sdc file.
@@ Appendix C More Results
C.1 Additional Analysis
In this section, we provide a detailed analysis of the layout results for the ariane133 benchmark, focusing on the observed relationship between wirelength, timing, area, and power. The AutoDMP method shows interesting yet seemingly contradictory outcomes—reduced wirelength with worse timing, smaller area, and lower power. To understand these results, we explore the factors that contribute to these trends and propose potential directions for placement algorithm improvements.
A key factor influencing timing in large-scale circuits is buffer insertion. In general, when the wirelength between two pins is large, signal transmission delay increases. Buffer insertion helps by breaking long interconnects into smaller segments, reducing the overall transmission delay. Furthermore, buffers enhance driving capacity by reducing the load on the signal driver, which helps lower signal delay and improve timing. However, fewer buffers result in smaller area and lower power but also lead to degraded timing performance. In the case of ariane133, the AutoDMP method inserts fewer buffers during timing repair, resulting in smaller overall area and power consumption. However, the lack of sufficient buffers leads to worse timing outcomes, as seen in the experimental data. This explains why, despite achieving minimized wirelength, the timing performance does not improve accordingly.
The relationship between wirelength and timing is often assumed to be directly proportional; however, the analysis of ariane133 demonstrates that this is not always the case. Although the total wirelength (as indicated by the HPWL) is reduced, the timing performance does not improve proportionally. In fact, congestion is lower and wirelength is minimized, but critical paths still suffer from significant delays. This suggests that global wirelength optimization does not always lead to improved timing, particularly for worst-case paths. Table 6 provides insights into this phenomenon by highlighting the inconsistencies between wirelength and timing slack on the worst paths. Specifically, the data reveals that longer wirelength paths do not always exhibit worse timing, indicating that the optimization strategy may not have sufficiently accounted for timing-sensitive paths.
To address these issues and improve overall placement quality, several optimization strategies can be explored. One potential solution to mitigate timing degradation is to apply weighted optimization that emphasizes critical paths. By giving more weight to timing-sensitive paths during wirelength reduction, it would be possible to achieve better alignment between wirelength optimization and timing improvement. Another approach is to adopt mixed-size placement techniques that optimize both macros and standard cells simultaneously. Since macro wirelength (macro HPWL) has a weaker correlation with backend performance metrics, this approach could lead to a more balanced placement outcome, particularly for designs where macro placement significantly impacts PPA. Additionally, machine learning models that predict PPA based on placement features could provide valuable insights during placement optimization. These models would allow the algorithm to dynamically adjust its strategy based on predicted PPA outcomes, leading to more informed trade-offs between PPA.
Table 6: Comparison of the worst path wirelength, delay, and slack for various methods on the ariane133 layout.
Method WireMask-EA SA MaskPlace ChiPFormer DREAMPlace AutoDMP OpenROAD
wirelength 3892 4107 3529 5761 4612 4318 4121
delay 4.36 4.57 4.25 4.49 4.54 4.92 4.44
slack -0.42 -0.54 -0.35 -0.55 -0.54 -0.98 -0.51
C.2 Raw Data
All the raw data from the experiment are in Tables 8-27.
@@ C.2 Raw Data
3 ariane136 The-OpenROAD-Project Ariane Core
4 bp The-OpenROAD-Project Full 64-bit RISC-V Core with Cache Coherence Directory
@@
7 CAN-BusTommydag A CAN bus Controller
8 DE2_CCD_edge suntodai Image processing
9 dft48Brendon Chetwynd, Kevin Bush, Kyle Ingols DFT design
10 FPGA-CANWangXuan95 A lightweight CAN bus controller
12 mor1kx openrisc (b) an OpenRISC processor IP core
13 or1200openrisc (a) OpenRISC 1200 implementation
14 OV7670_i2c AngeloJacobo Camera interface
@@
20 toygpu matt-kimball A simple GPU on a TinyFPGA BX
@@
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 1124169 5065453.5 0.226 6583143 0.369 -0.417 -329.353 1970 349228
SA 1683330 5187015.8 0.230 6699484 0.365 -0.512 -650.399 2317 347043
MaskPlace 4444289 6253554.3 0.265 7853892 0.373 -0.349 -244.936 1636 357322
ChiPFormer 1253799 5019138.5 0.226 6581086 0.370 -0.553 -860.952 2703 349005
DREAMPlace 1111023 4826654.0 0.214 6348638 0.367 -0.540 -690.266 2307 348180
AutoDMP 828592 4250870.4 0.192 5694373 0.350 -0.982 -1913.660 3310 344091
OpenROAD 2685856 5260791.9 0.235 6825071 0.359 -0.498 -596.718 2274 352706
@@
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 1181404 5278467.4 0.239 6945252 0.570 -1.726 -4268.610 3628 358740
SA 1693986 5608216.9 0.248 7208805 0.571 -1.620 -3917.960 3487 359532
MaskPlace 4586233 6544926.4 0.285 8245218 0.583 -1.508 -3677.480 3966 363704
ChiPFormer 1269027 5244085.2 0.237 6869186 0.566 -1.385 -3603.530 3609 358773
DREAMPlace 1067974 5202240.5 0.231 6831531 0.571 -1.358 -3269.220 4350 356706
AutoDMP 870515 4993936.2 0.228 6590972 0.562 -1.553 -4236.020 4431 356714
OpenROAD 3067334 5561870.0 0.246 7124942 0.546 -1.420 -3707.930 3473 360120
Table 10: The results of bp
@@
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 112560 7526021.6 0.437 10002159 0.254 -1.936 -21.812 326 490622
SA 151696 6927899.3 0.398 9113560 0.253 -1.965 -143.317 969 489850
MaskPlace 324368 7135790.7 0.407 9307133 0.253 -1.625 -44.217 478 489867
ChiPFormer 138943 6857073.6 0.392 8970666 0.250 -1.752 -502.977 2179 490305
DREAMPlace 161581 6989990.7 0.409 9347541 0.250 -2.108 -14.609 192 489693
AutoDMP 135090 6568910.6 0.376 8610145 0.247 -1.791 -254.786 1821 486445
OpenROAD 119123 6389902.6 0.365 8365647 0.247 -1.674 -219.229 1095 484494
Table 11: The results of bp_be
@@
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 66729 2678100.1 0.524 3574875 0.467 -2.142 -4093.970 5131 123966
SA 85984 2568796.6 0.557 3788559 0.459 -2.494 -5510.140 5381 124667
MaskPlace 114314 2548230.4 0.513 3522611 0.456 -2.439 -4459.870 5173 124938
ChiPFormer 62168 2525703.6 0.524 3572070 0.425 -2.169 -3541.820 5100 110904
DREAMPlace 87304 2533359.4 0.516 3518916 0.458 -2.163 -3648.020 6026 125221
AutoDMP 86263 2539005.8 0.468 3227167 0.453 -1.948 -3351.120 5045 125471
OpenROAD 76561 2370091.6 0.409 2835542 0.409 -1.862 -2343.290 5013 118786
Table 12: The results of bp_fe
@@
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 45985 1863651.1 0.510 2783740 0.309 -1.666 -777.420 2628 77648.300
SA 59872 1603778.9 0.404 2218668 0.310 -1.179 -1236.880 2990 81919.500
MaskPlace 64024 1540944.3 0.355 1965209 0.309 -1.318 -771.363 2582 77881.100
ChiPFormer 45754 1544281.0 0.375 2073376 0.299 -1.197 -1000.190 2714 72723.900
DREAMPlace 60279 1975681.6 0.508 2823861 0.294 -1.117 -473.261 1849 77427.800
AutoDMP 56107 1560612.6 0.365 2023847 0.312 -1.253 -1198.430 2626 81294.400
OpenROAD 53842 1468799.8 0.319 1770176 0.285 -1.092 -491.698 1870 75884.500
Table 13: The results of dft48
@@
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 335246296 1039236.4 0.114 1330932 0.239 -0.510 -439.378 2035 81353.700
SA 339003863 1815564.2 0.198 2305190 0.304 -0.882 -422.568 1546 84170.100
MaskPlace 1409226442 2117408.1 0.223 2552954 0.258 -0.381 -98.916 548 83185.400
ChiPFormer 431188529 1094111.0 0.119 1366989 0.242 -0.561 -235.601 974 80770.400
DREAMPlace 267700115 800165.5 0.112 1308094 0.234 -1.034 -801.257 2274 79407.400
AutoDMP 270017585 774418.4 0.091 1051008 0.232 -0.453 -125.948 797 80201.100
OpenROAD 752297935 1981412.2 0.198 2267002 0.257 -0.660 -393.111 1411 84300.200
Table 14: The results of or1200
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 102261063 1196931.6 0.035 1432301 0.049 -1.417 -1530.180 1754 66547.600
SA 113496323 1200696.8 0.034 1420640 0.059 -1.294 -1613.410 2674 66937.300
MaskPlace 208736767 1166784.1 0.033 1350820 0.059 -1.505 -2198.370 2684 66953.500
ChiPFormer 95035903 1089714.1 0.030 1262132 0.057 -1.208 -1415.960 2669 66256.900
DREAMPlace 153596940 1099937.2 0.031 1288746 0.057 -1.151 -1380.980 2670 65981.000
AutoDMP 122473700 1058807.4 0.033 1373410 0.058 -2.025 -3119.280 2760 65756.800
OpenROAD 187077195 1223792.9 0.034 1405193 0.060 -1.278 -1595.730 2677 67276.500
Table 15: The results of swerv_wrapper
Method Intermediate Metrics PPA Metrics
MacroHPWL $\downarrow$ HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
WireMask-EA 474447 3893655.8 0.414 4854661 0.666 -1.027 -873.506 1518 205627
SA 630919 3811960.2 0.368 4310596 0.648 -0.962 -854.737 1811 200280
MaskPlace 1270726 4193788.0 0.452 5286392 0.690 -1.251 -1306.460 1696 206481
ChiPFormer 437638 4060188.1 0.428 5019849 0.674 -1.189 -1282.240 2139 205775
DREAMPlace 865544 3575327.7 0.366 4525348 0.646 -1.061 -780.196 1608 203896
AutoDMP 450662 3213155.0 0.372 4354901 0.648 -1.094 -773.339 1529 200823
OpenROAD 666704 3177043.0 0.339 3997163 0.628 -1.127 -1163.820 1515 202853
Table 16: The results of 8051-master
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 142162.6 0.201 210127 0.075 -0.634 -16.460 35 29484.200
AutoDMP 131174.7 0.186 195536 0.074 -0.633 -15.643 35 29365.600
OpenROAD 141245.3 0.201 207737 0.073 -0.603 -16.664 41 29461.400
Table 17: The results of CAN-Bus
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 3916.8 0.093 5683 0.005 -0.074 -0.074 1 1675.270
AutoDMP 4210.8 0.097 6024 0.005 -0.074 -0.074 1 1668.880
OpenROAD 3955.1 0.092 5623 0.005 -0.079 -0.079 1 1665.430
Table 18: The results of DE2_CCD_edge
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 16556.5 0.123 23453 0.201 -0.924 -184.592 736 5797.470
AutoDMP 22203.0 0.160 30745 0.211 -0.930 -200.125 710 5812.630
OpenROAD 16689.0 0.120 22746 0.199 -0.919 -165.009 609 5758.100
Table 19: The results of FPGA-CAN
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
AutoDMP 1437916.7 0.178 2393640 0.745 -0.892 -108.201 3864 379675
OpenROAD 1483498.8 0.180 2394544 0.742 -0.404 -30.331 1843 379193
Table 20: The results of OV7670_i2c
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 733113.0 0.215 1164763 0.625 -0.579 -82.237 3154 157496
AutoDMP 701777.0 0.201 1091553 0.612 -0.561 -23.030 1053 155624
OpenROAD 723012.1 0.208 1127120 0.617 -0.466 -30.612 2327 156210
Table 21: The results of iot_shield
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 5014.3 0.118 7346 0.006 -0.167 -2.457 27 1712.770
AutoDMP 5976.2 0.137 8653 0.006 -0.152 -2.367 29 1730.600
OpenROAD 5067.3 0.119 7349 0.006 -0.152 -2.186 20 1719.420
Table 22: The results of mor1kx
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
@@
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
@@
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
@@
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
@@
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
@@
Table 27: The results of toygpu
Method Intermediate Metrics PPA Metrics
HPWL $\downarrow$ Congestion $\downarrow$ Wirelength $\downarrow$ Power $\downarrow$ WNS $\uparrow$ TNS $\uparrow$ NVP $\downarrow$ Area $\downarrow$
DREAMPlace 4347733.8 0.198 6784639 1.101 -1.503 -131.661 101 979293
AutoDMP 4444860.0 0.194 6595166 1.109 -2.418 -201.797 107 967780
OpenROAD 4657703.6 0.212 7185236 1.108 -1.294 -103.668 101 981911